OpenStudy (anonymous):

what are the zeros of a cotangent function? and does amplitude affect cotangent functions?

OpenStudy (anonymous):

\[\cot(x)=\frac{\cos(x)}{\sin(x)}\] so zeros of cotangent same as zeros of cosine

OpenStudy (anonymous):

that was wrong. sorry

OpenStudy (anonymous):

zeros are \[\frac{\pi}{2}, \frac{3\pi}{2}...\]

OpenStudy (anonymous):

so cot theda=o when cos theda=0

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

odd multiples of pi over 2 \[\frac{(2k-1)\pi}{2}, k \in Z\]